# It Takes a Mirror to Find Flaws: Uncovering Biases with Reflective Large Language Models (Supplementary Materials)

### v1

@@
Appendix B: Experiment #2, Debate #2
Table 1 provides an overview of the discussion surrounding the proposition: “Quantum entanglement has the potential to significantly transform the fields of computing and artificial intelligence.” This debate is particularly pertinent given the escalating excitement and, occasionally, the excessive hype associated with quantum computing’s capabilities and future impact. By engaging in this debate, we aim to critically evaluate and place the potential of quantum entanglement in computing and artificial intelligence into an accurate and measured perspective.
The structure of the debate is delineated through the participation of two distinct agents. Agent 1 assumes the role of the proponent, advocating for the transformative power of quantum entanglement in revolutionizing computing and artificial intelligence. In contrast, Agent 2 adopts a skeptical stance, challenging the optimistic projections and questioning the current state and near-future possibilities offered by quantum technologies.
Agent #1 Supporting Arguments Agent #2 Opposing Arguments
1. My key arguments center around several significant aspects and developments. These include empirical evidence proving quantum entanglement through experiments like Bell’s Theorem, the immense potential and emerging breakthroughs of quantum computing (e.g., Google’s Sycamore processor), the possible transformative effects on AI, the inception of revolutionary quantum algorithms, the prospects of quantum networks, quantum enhancement of data security through unbreakable encryption, and the potential for birthing new quantum-based technologies. 2. While recognizing the concerns raised by Agent2 about experimental limitations, the challenges in erecting scalable, efficient quantum computers, hurdles in quantum error correction and AI integration, the limitations of existing quantum algorithms, infrastructure and training obstacles, and skepticism about the extent of the quantum revolution, I have drawn attention to the ongoing efforts and substantial progress in these areas. 3. For instance, experimental entanglement has been successfully demonstrated despite their controlled settings, large strides are being made in the scalability and efficiency of quantum computers, error correction methodologies are being developed and refined, Quantum Machine Learning heralds a promising intersection of quantum computing and AI, and new quantum algorithms are an active area of research. 4. A marked difference between Agent2 and me lies in our outlook towards the future of quantum entanglement. While both of us acknowledge the challenges, I express robust optimism. 1. The central arguments for my stance include experimental uncertainties, quantum computing hurdles, a high quantum error rate, the challenges with incorporating quantum computing into AI, the early stage of quantum algorithms, and the demands on infrastructure and human resources. 2. These counterpoints refute many of Agent1’s affirmations. Despite the scientific proof of quantum entanglement, it’s essential to understand that these proofs come from controlled lab environments, often wrestling with loopholes and scalability issues. While Google’s demonstration of quantum supremacy showcases the potential, it does not invalidate the fact that quantum computers, as of now, are delicate, error-prone and largely inefficient compared to classical ones. 3. Concerning AI, while theoretically, quantum computing could indeed provide benefits, current AI systems are fully entrenched in classical computing. Moreover, the inherent instability and error-proneness of quantum computing currently make them less suitable for AI tasks that rely on vast amounts of real-world data. 4. Significant differences exist between our positions. While Agent1 seems eager to champion the potential of quantum technology, I lean towards a more cautious and practical assessment, highlighting the significant hurdles we currently face in making quantum computing and AI a reliable, efficient reality.
In summary, our debate has provided a comprehensive overview of the fascinating realm of quantum entanglement, its potential impact on computing and AI, and the associated challenges. It is clear that we are at the cusp of a quantum revolution, albeit faced with substantial hurdles. While the speed and scope of this change remain open questions, the sheer potential of quantum entanglement makes it an exciting field to watch as we venture into the quantum era. In summary, the future of quantum entanglement in computing and AI is still uncertain with numerous unresolved issues. These include developing stable, practical quantum computers, creating quantum algorithms applicable to a broad range of problems, and successfully integrating quantum computing into classical systems and AI. It is crucial not to overlook the vast scientific and practical challenges that precede it.
Table 1: Debate subject: Quantum entanglement holds tremendous potential to revolutionize computing and advance artificial intelligence?
Appendix C: Training Dataset
The training dataset used in Section 5 were submitted on Open Review under the Data section. The link provided in reference SocraSynthBiasesDataSet can also access the dataset.
Appendix Z: Additional Supplementary Materials
For additional experimental results, datasets, and debate interactions, please refer to our supplementary materials ACLSupplementary.
Appendix B: Experiment #1 Justifications of 31 Biased Articles
In Section LABEL:sec:exp1, we note that EVINCE processed 31 news articles to assess their neutrality. In addition to the final decision, we detail the justifications EVINCE provides at the debate’s end. These justifications are documented in four tables: Tables 2, 3, 4, and 5.
# Rating S,D,R EVINCE Justifications
@@
D1 5,5,5 The article discusses the controversy surrounding the IRS’s alleged targeting of conservative groups, focusing on Rep. Darrell Issa’s stance against releasing full transcripts from IRS interviews. Issa argues that full disclosure would compromise the investigation, a position criticized by Rep. Elijah Cummings, who advocates for transparency. The debate highlights political tensions, with implications of bias in the handling of the investigation. The article portrays a complex interplay of political, ideological, and possibly historical biases, reflecting the contentious nature of the IRS scandal and its investigation.
@@
Table 2: The First 9 of 16 Democrat Scandals News Articles Rated by EVINCE and its Justifications. The rating column starts with EVINCE’s rating, the Democrat rater (in blue), and then Republican rater (in red).
@@
D278 4,4,3 The article discusses a House committee investigation into whether Attorney General Eric Holder misled Congress regarding the Department of Justice’s pursuit of journalists’ personal records. It highlights Holder’s previous testimony claiming no involvement in prosecuting press for material disclosure, contrasted with the DOJ’s actions against Fox News reporter James Rosen, who was labeled a potential "co-conspirator" but not charged…. The article’s tone suggests potential bias by emphasizing allegations that Eric Holder may have misled Congress, primarily focusing on criticisms from Republican figures and conservative critics. It uses specific incidents to question Holder’s integrity and the Justice Department’s actions under his leadership, possibly framing the narrative to highlight governmental overreach and lack of transparency.
@@
Table 3: The Last 7 of 16 Democrat Scandals News Articles Rated by EVINCE and its Justifications. The rating column starts with EVINCE’s rating, the Democrat rater (in blue), and then Republican rater (in red).
# Rating S,D,R EVINCE Justifications
@@
R7 3,3,3 The article reports on allegations by Senator Mitch McConnell that his campaign headquarters were wiretapped, with the FBI investigating these claims. A recording of McConnell’s team discussing potential attacks on Ashley Judd, who was considering running against him, was released by Mother Jones. McConnell accused the political left of this action, describing it as a "Nixonian move." The recording included discussions on various strategies to undermine potential opponents, highlighting a focus on Judd’s personal struggles and political views. The controversy has prompted responses from both Republican and Democratic officials, reflecting the tense political atmosphere.
@@
R69 4,4,4 The report discusses how young Republicans are seeking a different message for elections, emphasizing a departure from divisive social issues and a focus on fiscal responsibility, national defense, and energy advancement. Selection Bias: The article primarily focuses on young Republicans who are seeking a different message for the party. It doesn’t provide as much insight into young Republicans who may still align with traditional conservative values, which could create a slight bias toward the viewpoints of those seeking change. Language Bias: Certain language choices, such as describing divisive social issues as "anti-abortion, anti-gay, and anti-environment stances," may reflect a bias toward more progressive viewpoints on these issues. A more neutral description might be "positions on abortion, same-sex marriage, and environmental policy." Source Bias: The perspectives provided in the article are mainly from young Republicans themselves. While including these voices is essential, the article could benefit from additional perspectives from political analysts or experts to provide more context and balance.
@@
Table 4: The First 8 of 15 Republican Scandals News Articles Rated by EVINCE and its Justifications. The rating column starts with EVINCE’s rating, the Democrat rater (in blue), and then Republican rater (in red).
# Rating S,D,R EVINCE Justifications
@@
R221 3,3,4 "Hurricane Christie" presents Governor Chris Christie’s critique of House Republicans in a manner that emphasizes party conflict and personal betrayal. The dramatic framing, choice of language, and focus on internal discord may introduce bias by portraying Christie’s actions in a specific light and emphasizing the divide within the Republican Party. The article’s approach to presenting these events can influence readers’ perceptions, potentially leading them to see the situation through a lens of heightened drama and internal strife.
@@
R280 5,5,3 The article from Fox News by Jay Sekulow, titled "Obama’s fingerprints all over IRS Tea Party scandal," presents a viewpoint that directly implicates President Obama in the IRS scandal involving the targeting of conservative groups. The author argues that the scandal was not only known but encouraged by senior IRS officials, Congressional Democrats, the White House, and further fueled by the mainstream media. To assess the article for bias, let’s evaluate it against various criteria: The article "Obama’s fingerprints all over IRS Tea Party scandal" demonstrates clear political and ideological biases, with a narrative constructed to directly implicate President Obama in the IRS targeting scandal. By selectively quoting Obama and drawing connections to actions by the IRS, the article aims to present a cohesive narrative that places responsibility for the scandal on the president. This framing serves to reinforce the viewpoint of those who see the actions as politically motivated and indicative of broader issues of governance and accountability under the Obama administration. The choice of language, historical comparisons, and the leveraging of the author’s and platform’s ideological stances contribute to a biased presentation of the events surrounding the IRS scandal.
Table 5: The Last 7 of 15 Republican Scandals News Articles Rated by EVINCE and its Justifications. The rating column starts with EVINCE’s rating, the Democrat rater (in blue), and then Republican rater (in red).
Appendix C: Experiment Dataset

### v2

@@
Appendix C: Experiment #1 Justifications of 31 Biased Articles
[…]
@@
D1 4,4,4 The article discusses the controversy surrounding the IRS’s alleged targeting of conservative groups, focusing on Rep. Darrell Issa’s stance against releasing full transcripts from IRS interviews. Issa argues that full disclosure would compromise the investigation, a position criticized by Rep. Elijah Cummings, who advocates for transparency. The debate highlights political tensions, with implications of bias in the handling of the investigation. The article portrays a complex interplay of political, ideological, and possibly historical biases, reflecting the contentious nature of the IRS scandal and its investigation.
@@
[…]
@@
D278 4,4,3 The article discusses a House committee investigation into whether Attorney General Eric Holder misled Congress regarding the Department of Justice’s pursuit of journalists’ personal records. It highlights Holder’s previous testimony claiming no involvement in prosecuting press for material disclosure, contrasted with the DOJ’s actions against Fox News reporter James Rosen, who was labeled a potential “co-conspirator” but not charged…. The article’s tone suggests potential bias by emphasizing allegations that Eric Holder may have misled Congress, primarily focusing on criticisms from Republican figures and conservative critics. It uses specific incidents to question Holder’s integrity and the Justice Department’s actions under his leadership, possibly framing the narrative to highlight governmental overreach and lack of transparency.
@@
[…]
@@
R7 3,3,3 The article reports on allegations by Senator Mitch McConnell that his campaign headquarters were wiretapped, with the FBI investigating these claims. A recording of McConnell’s team discussing potential attacks on Ashley Judd, who was considering running against him, was released by Mother Jones. McConnell accused the political left of this action, describing it as a “Nixonian move.” The recording included discussions on various strategies to undermine potential opponents, highlighting a focus on Judd’s personal struggles and political views. The controversy has prompted responses from both Republican and Democratic officials, reflecting the tense political atmosphere.
@@
R69 4,4,4 The report discusses how young Republicans are seeking a different message for elections, emphasizing a departure from divisive social issues and a focus on fiscal responsibility, national defense, and energy advancement. Selection Bias: The article primarily focuses on young Republicans who are seeking a different message for the party. It doesn’t provide as much insight into young Republicans who may still align with traditional conservative values, which could create a slight bias toward the viewpoints of those seeking change. Language Bias: Certain language choices, such as describing divisive social issues as “anti-abortion, anti-gay, and anti-environment stances,” may reflect a bias toward more progressive viewpoints on these issues. A more neutral description might be “positions on abortion, same-sex marriage, and environmental policy.” Source Bias: The perspectives provided in the article are mainly from young Republicans themselves. While including these voices is essential, the article could benefit from additional perspectives from political analysts or experts to provide more context and balance.
@@
[…]
@@
R221 3,3,4 “Hurricane Christie” presents Governor Chris Christie’s critique of House Republicans in a manner that emphasizes party conflict and personal betrayal. The dramatic framing, choice of language, and focus on internal discord may introduce bias by portraying Christie’s actions in a specific light and emphasizing the divide within the Republican Party. The article’s approach to presenting these events can influence readers’ perceptions, potentially leading them to see the situation through a lens of heightened drama and internal strife.
@@
R280 5,5,3 The article from Fox News by Jay Sekulow, titled “Obama’s fingerprints all over IRS Tea Party scandal,” presents a viewpoint that directly implicates President Obama in the IRS scandal involving the targeting of conservative groups. The author argues that the scandal was not only known but encouraged by senior IRS officials, Congressional Democrats, the White House, and further fueled by the mainstream media. To assess the article for bias, let’s evaluate it against various criteria: The article “Obama’s fingerprints all over IRS Tea Party scandal” demonstrates clear political and ideological biases, with a narrative constructed to directly implicate President Obama in the IRS targeting scandal. By selectively quoting Obama and drawing connections to actions by the IRS, the article aims to present a cohesive narrative that places responsibility for the scandal on the president. This framing serves to reinforce the viewpoint of those who see the actions as politically motivated and indicative of broader issues of governance and accountability under the Obama administration. The choice of language, historical comparisons, and the leveraging of the author’s and platform’s ideological stances contribute to a biased presentation of the events surrounding the IRS scandal.
[…]
Appendix D: Experiment Dataset